A rectangular parking lot is 50 ft longer than it is wide. Determine the dimensions of the parking lot if it measures 250 ft diagonally.
Answer by checkley77(12844)

x^2+(x+50)^2=250^2
x^2+x^2+100x+2,500=62,500
2x^2+100x+2,500-62,500=0
2x^2+100x-60,000=0
2(x^2+50x-30,000)=0
2(x-150)(x+200)=0
x-1150=0
x=150 for one side.
150+50=200 for the other side.
Proof:
150^2+200^2=250^2
22,500+40,000=62,500
62,500=62,500